A guitar string is 60 cm long. A pluck on the string sends it into its third harmonic. If the frequency of this harmonic is 880 Hz and the tension is 1200 N, what is the linear density of the string in kg/m?
Answer #1

f = 1/2l(sqrt(T/m))

f3 = 880 = 3f

f = 293.3 Hz

T/m = 4l^2f^2

m = 1200/4*0.6^2*293.3^2

m = 0.00969 kg/m